OpenStudy (anonymous):

Use the Rational Zeros Theorem to write a list of all possible rational zeros of the function. f(x) = -2x4 + 4x3 + 3x2 + 18

terenzreignz (terenzreignz):

two key numbers here: the leading coefficient, and the constant. What are they?

OpenStudy (anonymous):

i have no clue.

terenzreignz (terenzreignz):

Leading coefficient is the number multiplied to the x with the *highest exponent* while the constant, is, well, the constant (the number with no x attached to it) So... what are the leading coefficient and the constant (respectively)?

OpenStudy (anonymous):

2x^4 and 18?

terenzreignz (terenzreignz):

drop the x. And it's actually -2.

terenzreignz (terenzreignz):

Now, here's what needs to be done... I want you to list down all the factors of 2 and also all the factors of 18.

OpenStudy (anonymous):

2 2,1 18 1,18 2,9 3,6

terenzreignz (terenzreignz):

There really is no need to pair them up :) But anyway, well done :P So, to get the *possible* rational root, just pair them up as fractions, with the factors of 18 in the numerator and factors of 2 in the denominator. In other words, the possible rational roots take the form \[\Large \pm\frac{q}p\] Where q is a factor of 18 and p is a factor of 2. Don't be redundant, though... there is no need to include, say, both \[\Large \pm\frac{6}2\color{orange}{=3} \text{ and} \ \pm \frac{3}{1} \color{orange}{=3}\] Since they are obviously the same

OpenStudy (anonymous):

im confused..

OpenStudy (anonymous):

so like 9/2 2/1 ?

terenzreignz (terenzreignz):

Just list down all possible pairings of a factor of 18 and a factor of 2. For example, 6 (factor of 18) and 2 (factor of 2) And just divide them like so: (don't forget the plus-minus) \[\Large \pm\frac63= \pm 2\] For example, say the leading coefficient was 2 and the constant was 6. factors of 2 = {1,2} factors of 6 = {1,2,3,6} then, the possible rational roots would be all possible divisions between factors of 6 and 2. \[\Large PRR = \left\{\pm\frac61 , \pm\frac62,\pm\frac21, \pm\frac32,\pm\frac22, \pm\frac12\right\}\]\[\Large PRR = \left\{\pm 6,\pm 3, \pm 2\pm \frac32,\pm1,\pm\frac12\right\}\] I didn't include 3/1 because it's the same as 6/2 I didn't include 1/1 because it's the same as 2/2 But that's all of it :)

terenzreignz (terenzreignz):

Now, your constant is 18 (not 6) so expect more possible rational roots... but the essence remains the same :D Good luck ^_^

OpenStudy (anonymous):

ohhh thank you so much!

terenzreignz (terenzreignz):

No problem ^_^

terenzreignz (terenzreignz):

Post your answer if you can.